The larger of two numbers is three less than four times the smaller number. If eight times the larger number is subtracted from 20, the result is -20. Find each of the numbers.... Please help me I don't understand how to do these problems!!!!!!!
Answer by lwsshak3(11628)

The larger of two numbers is three less than four times the smaller number. If eight times the larger number is subtracted from 20, the result is -20. Find each of the numbers.
**
let x=smaller number
4x-3=larger number (three less than four times the smaller number)
20-8(4x-3)=-20 (eight times the larger number is subtracted from 20, the result is -20
20-32x+24=-20
32x=20+24+20=64
32x=64
x=2
4x-3=8-3=5
ans:
smaller number=2
larger number=5